Session: Jun 2013 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roma at Mon Jun 03 05:31:04 2013) Success:
All scan recorded. Sky clear.

Noto (g.nicotra at Mon Jun 03 05:33:49 2013) Success:

Yebes (Daniel Valero at Mon Jun 03 06:23:45 2013) Success:
No known problems

Effelsberg (Georgi at Mon Jun 03 07:01:20 2013) Success:

Torun (K. Borkowski at Mon Jun 03 07:52:37 2013) Success:
No known problems.

Westerbork (Geert Kuper at Mon Jun 03 10:01:09 2013) Success:
!Note: to avoid a tadumax problem we steered DSB but recorded USB only as requested

Onsala (Michael Lindqvist at Tue Jun 04 09:45:35 2013) Failure:
Could not observe due to problems with the 220 V to the antenna control.